### v3

# Efficient IoT Intrusion Detection with an Improved Attention-Based CNN-BiLSTM Architecture

Abstract
The ever-increasing security vulnerabilities in the Internet-of-Things (IoT) systems require improved threat detection approaches. This paper presents a compact and efficient approach to detect botnet attacks by employing an integrated approach that consists of traffic pattern analysis, temporal support learning, and focused feature extraction. The proposed attention-based model benefits from a hybrid CNN-BiLSTM architecture and achieves 99% classification accuracy in detecting botnet attacks utilizing the N-BaIoT dataset, while maintaining high precision and recall across various scenarios. The proposed model’s performance is further validated by key parameters, such as Mathews Correlation Coefficient and Cohen’s kappa Correlation Coefficient. The close-to-ideal results for these parameters demonstrate the proposed model’s ability to detect botnet attacks accurately and efficiently in practical settings and on unseen data. The proposed model proved to be a powerful defence mechanism for IoT networks to face emerging security challenges.
Index Terms:
Botnets, intrusion detection, BiLSTM, IoT, attention mechanism.
I Introduction
The application of Artificial Intelligence (AI) in Internet-of-Things (IoT) has resulted in innovative and sophisticated networked systems. Such systems, comprising of a number of devices ranging from simple sensors to advanced machinery, influence a number of applications, e.g., smart cities, homes, transportation, manufacturing, and agriculture [1]. IoT devices have become deeply embedded in everyday life. However,this extensive adoption brings significant security risks. Due to their diverse nature and inherent vulnerabilities, IoT devices are particularly prone to malicious activities or cyber attacks such as malware attacks. As the use of IoT devices continues to grow, concerns over security and privacy are rising, given the increasing exposure of sensitive data online. To reduce these cyber attacks, continuous network monitoring is crucial. Analysing IoT network traffic can enhance intrusion detection mechanisms within the IoT ecosystem, ultimately strengthening the cybersecurity of networks [2]. Intrusion detection systems serve as essential supplementary security measures, operating in conjunction with other defences to reduce potential threats. Deep Learning, a subset of machine learning, is capable of extracting relevant information from massive and high-speed network traffic generated by various interconnected IoT devices [3].
This paper introduces a hybrid deep learning security framework that integrates an attention mechanism with 1D convolutional neural networks (CNN) and Bidirectional Long Short-Term Memory (BiLSTM) architectures. The framework demonstrates robust intrusion detection capabilities using normalized data and measuring performance using the N-BaIoT dataset across various performance and reliability metrics, taking into account the specific challenges of the datasets, including device behaviours and class imbalance [4]. This study presents the following key contributions:
  1. 1.
    A hybrid attention-based deep learning model that integrates 1D-CNN + BiLSTM architectures to efficiently and accurately detect botnet attacks.
  2. 2.
    The proposed model demonstrates a reduced computational overhead without compromising the detection accuracy.
  3. 3.
    The proposed model has been compared with the latest deep learning approaches demonstrating its superior performance.
II Literature Review
The increase in the number of resource-constrained IoT devices has made their security more challenging, as novel security risks emerge. Recently, researchers have been developing a lot of cryptography techniques [5, 6] and artificial intelligence-based IDS to protect IoT systems. A number of machine learning methods can be found that utilize N-BaIoT data to address these security concerns, for example, in [7], the authors presented a collaborative learning model for an improved IDS intended to secure IoT industry 4.0. Their methodology, which merge cooperative learning and Deep Neural Networks (DNN) approaches, scored well, obtaining an 0.92% as F1-Score and a 95.06% of accuracy. Another study [8], utilising several neural network structures demonstrated that CNN, RNN and LSTM outperformed in botnet identification with an accuracy of 91%, 41% and 62% respectively. Moreover, the authors in [9] developed a CNN-based deep learning model for botnet detection using the N-BaIoT dataset, improving scalability in diverse environments. The model showed superior accuracy, precision, and recall, but struggled with UDP-based attacks. The research emphasizes the effectiveness of CNNs in IoT security. Similarly, a novel CNN-LSTM model has been introduced in [10], which detects BASHLITE and Mirai security attacks, achieving an accuracy of 90.88% and precision and recall rates of 93.04% and 91.91% respectively. This work marks a significant improvement in attack detection capabilities. These studies demonstrate the advancement and efficacy of several algorithms in addressing the challenges connected with IoT botnet detection.
III Preliminaries
The proposed model integrates 1D-CNN, BiLSTM and an attention mechanism to create a robust architecture for sequential data analysis. This hybrid approach capitalizes on CNN’s ability to detect local patterns and spatial relationships, BiLSTM’s capacity to capture long-range dependencies, and the attention mechanisms focus on crucial input elements.
III-A 1D-CNNs
1D-CNNs have shown particularly effective for classification methods, adopting a multi-layered design that consists of five different layers for processing.The Convolutional layers serve a substantial part in determining feature extraction and parameter optimization during back propagation. These structures make use of dedicated filters, incorporating pooling techniques to prevent over-fitting. The data is then turned into class representations using flattening and connected layers. The modular design provides structural flexibility to meet a variety of categorization requirements. The mathematical equation[11] of the 1D – CNN are as follows:
$y_{i}^{l}=\sigma\left(\sum\limits_{i=1}^{N_{l-1}}\text{conv1D}(w_{i,j}^{l},x_{i}^{l-1})+b_{j}^{l}\right)$ (1)
III-B Bidirectional Long Short-Term Memory (Bi-LSTM)
Bi-LSTM analyses sequences bidirectionally to detect connections and temporal patterns. Its nature of two-way analysis improves its ability to recognize complicated sequential patterns.
Figure 1: The BiLSTM Block [12].
As this architecture employs two hidden layers: backward and forward (in Fig:1) The forward layers processes sequence from start to end, whereas the backward layer processes it in reverse way. After it combines the results from both layers as final output.Some important equations of that particular model are as follows:
$h_{t}^{f}=\tanh(W_{xh}^{f}x_{t}+W_{hh}^{f}h_{t-1}^{f}+b_{h}^{f})$ (2)
$h_{t}^{b}=\tanh(W_{xh}^{b}x_{t}+W_{hh}^{b}h_{t+1}^{b}+b_{h}^{b})$ (3)
$y_{t}=W_{hy}^{f}h_{t}^{f}+W_{hy}^{b}h_{t}^{b}+b_{y}$ (4)
III-C Attention Mechanism
The method utilizes dot product attention based on key, value and query connections.The key matrix, which originates from the value matrix, interacts with query vectors to present importance weights. These weights then lead process of aggregating value vectors.
$\mathbf{K}=\tanh(\mathbf{VW}^{a})$ (5)
$\mathbf{d}=\text{softmax}(\mathbf{qK}^{T}),\quad\mathbf{a}=\mathbf{dV}$ (6)
Figure 2: Attention Mechanism [13].
IV The Proposed Deep Learning Model
IV-A Hybrid model architecture
In this work, we used N-BaIoT dataset, which consists of network traffic data from five categories of IoT devices: security cameras, webcams, doorbells,thermostats, baby monitors. This includes both ordinary traffic and traffic generated by several botnet assaults, namely Mirai and Bashlite (gafgyt). We implemented a model approach to correctly classify network data as normal or pertaining to one of nine different attack types. To improve performance this neural network architecture consists of three components one dimensional convolutional, bidirectional long short-term memory (Bi-LSTM), and attention process. The model’s framework starts with a 1D convolutional with 128 filters, a kernel size of 5, and ReLU activation. This is preceded by batch normalization, maximum pooling and a 0.3% dropout rate. Two more convolutional layers are added, one with 256 filters and the second is with 128 filters with a kernel size of three, maintaining an identical pattern. Temporal features are eventually examined by a Bi-LSTM layer of 128 units, which is supported by an attention mechanism. The structure continues with two dense layers of 256 and 128 units each, all with 0.4 rate of dropout and with ReLU activation. The final layer yields a 10 class softmax output. The optimization process Adam with learning rate of 0.001 is used in the model.
Figure 3: The Proposed Model Architecture
V Experimental Setup
To validate our model, we used the N-BaIoT dataset, allocating 20% for testing and 80% for training purposes. The model was trained on an NVIDIA T4 GPU, and the total training time was approximately 1 hour. During inference, the model achieves an average inference time of around 50 milliseconds per sample. The model’s training setup included 2 dropout layers (0.3, 0.4), and the Adam optimizer, used sparse categorical cross-entropy with softmax activation to provide multi-class classification output.
TABLE I: Data Attack Schema
Botnet Attack Type Doorbell Thermostat Baby Monitor Webcam Security Camera 1 Security Camera 2
Benign 88648 13113 175240 52150 160688 6613
Bashlite COMBO 112732 53012 58152 58669 118910 113681
JUNK 58865 30312 28349 28305 59966 55992
SCAN 57969 27494 27859 27698 57694 56397
TCP 193677 95021 92581 97783 193897 186891
UDP 209807 104791 105782 110617 208669 206700
Mirai ACK 102195 113285 91123 0 118551 218667
SCAN 107685 43192 103621 0 193877 89604
SYN 122573 116807 118128 0 127597 248194
UDP 237665 151481 217034 0 314856 308963
Plain UDP 81982 87368 80808 0 110466 162680
Total 1373798 835876 1098677 375222 1665151 1713882
V-A Dataset
The N-BaIoT dataset [14]incorporates 7,062,606 network traffic data patterns from several smart devices to identify attacks like botnet. It comprises of 115 features determined across many time frames, involving both valid and malicious activities on the network. It includes traffic data from nine business IoT devices infected with malware either Mirai or BASHLITE,conveying 10 various attack types. The dataset has 10 classes of attacks with 1 class as benign.
Table 1. The above illustrates the schema and the number of rows that are present to apply our models.
V-A1 Data Gathering and Preprocessing
  • Data Gathering: Normal traffic data was collected immediately after the installation of the IoT devices. Subsequently, the devices were infected with different botnet attacks, and the resulting traffic was captured using Wireshark.
  • Feature Extraction: From the captured traffic, 115 features were extracted, encompassing statistical information over various time intervals ranging from 100 milliseconds to one minute.
  • Normalization: Given that the dataset consists solely of numeric features, we applied the Min-Max scaler to normalize the data. This scaling technique ensures that all features are within the range [0, 1], facilitating the training process and improving model performance.
  • Class Balancing: To address class imbalance, we ensured that each class (benign and the 9 types of botnet attacks out of 10) had an equal number of samples. Specifically, we used 10,000 samples per class, resulting in a balanced dataset that mitigates bias and enhances the robustness of our model.
V-B Evaluation Parameters
V-B1 Performance Parameters
To analyze the performance of the architecture, we used standard metrics resulting from the confusion matrix. These indicators comprise of accuracy, recall, precision, and F1-score all these provides a dept insight into the model’s classification capabilities. Four key elements are involved in confusion matrix: False Positives (FP)—Normal instances mistakenly classified as attacks, False Negatives (FN)—Attack instances mistakenly classified as normal, True Positives (TP)—Correctly identified attack instances, True Negatives (TN)—Correctly identified normal instances. Using these elements, we calculate the following indicators:
1. Accuracy: The percentage of accurate predictions (which involves true positives and true negatives) across all instances examined.
$\text{Accuracy}=\frac{TP+TN}{TP+FP+TN+FN}$ (7)
2. Recall: The output of accurately determined positive instances to the total number of positive instances.
$\text{Recall}=\frac{TP}{TP+FN}$ (8)
3. Precision: The proportion of precisely detected positive cases to all anticipated positive instances.
$\text{Precision}=\frac{TP}{TP+FP}$ (9)
4. F1- Score: The harmonic mean of recall as well as precision, presents a fair evaluation of the performance of respective model.
$F1\text{-Score}=2\times\frac{Precision\times Recall}{Precision+Recall}$ (10)
V-C Reliability Parameters
The evaluation procedure includes two advanced sophisticated statistical tools: Kappa coefficient and Matthew’s correlation. These additional measures improve the procedure by analyzing relationships between classes and considering factors of random agreement, providing a more thorough insight of classifier performance.
V-C1 Cohen’s Kappa Co-efficient
It is a statistic used for evaluating classification ability that examines the concordance between actual and anticipated results while incorporating agreements that may arise by chance [15]. This coefficient operates on a scale of negative one to positive one, with positive one indicating ideal agreement, zero indicating chance acceptance and negative one signifying contradiction. This technique provides a more detailed evaluation than traditional measures.
V-C2 MCC
It provides a sophisticated measure of classification quality, particularly useful for datasets with unbalanced class distribution [16]. This complete measure uses all elements of confusion matrix – TP, FP,TN, and FN, to calculate a score ranging from negative to positive one. A flawless anticipated accuracy is represented by a score of +1, 0 yielding a random guessing score, and -1 represents entirely opposite predictions. As a result, MCC is especially useful for evaluating classifier performance when working with unbalanced datasets.
VI Results and Analysis
VI-A Training Results
We appraise the results of the hybrid approach on the basis of recall, accuracy, precision, F1- score and correlation coefficients. The training results are shown in Figure 5.
Figure 4: Accuracy Curve
Figure 5: Loss Curve
  • Loss: The decreasing trend in the loss curves implies an ongoing reduction in the error rate of the architecture on the training dataset as epochs rise up. This pattern indicates that the model is actively improving and learning its predictive skills.
  • Accuracy: The accuracy curves rise correspondingly, showing that the proposed model’s predictions are growing accurately.
  • Stability Near 30th Epoch: The model’s accuracy stabilizes around the 30th Epoch, suggesting it has learned, effectively identify underlying trends in the data and beyond that it will start to overfit.
VI-B Confusion Matrix and Classification Report
Figure 6: Confusion Matrix
The confusion matrix is utilized to evaluate the classifiers performance comprehensively. When most of the samples in the matrix are concentrated along the diagonal, as illustrated in Figure 5, the model can accurately differentiate between attack samples and normal samples.
  • Class-Specific Metrics:
  • mirai_udp: High precision (1.00) but lower recall (0.93), suggesting the model is very accurate when it predicts this class, but it misses some instances.
  • mirai_ack: Lower precision (0.93) but perfect recall (1.00), meaning the model identifies all instances but has some false positives.
  • Remaining classes : High scores across all metrics, indicating strong performance with minor variations.
TABLE II: Classification report of each class
Class Precision Recall F1-Score Support Accuracy
benign 1.00 0.996 0.998 2034 0.997
gafgyt_combo 0.995 0.964 0.979 1989 0.964
gafgyt_junk 0.966 0.995 0.980 2010 0.996
gafgyt_scan 0.998 0.999 0.999 1988 0.999
gafgyt_udp 0.997 0.999 0.998 1932 0.928
mirai_ack 0.929 0.999 0.963 1976 0.999
mirai_scan 0.998 1.00 0.999 2026 1.00
mirai_syn 1.00 0.999 0.999 1950 0.998
mirai_udp 0.999 0.928 0.962 1991 0.928
mirai_udpplain 0.999 0.997 0.998 2104 0.997
accuracy 0.988 0.988 0.988 0.987
Macro-Average 0.988 0.988 0.988 20000
Weighted-Average 0.988 0.988 0.988 20000
VI-C Reliability Parameters
VI-C1 Kappa Coefficient
It assesses the agreement between actual and anticipated classifications after controlling for chance agreement. A kappa score of 0.985% indicates practically perfect agreement
VI-C2 Mathews Correlation Coefficient
It measures the level of binary classifications. It suggests 0.986% a high connection between expected and true classifications.
TABLE III: Cohen’s Kappa and Matthews Correlation Coefficient (MCC) for each class
Class Cohen’s Kappa Matthews
benign 0.998 0.998
gafgyt_combo 0.976 0.976
gafgyt_junk 0.978 0.978
gafgyt_scan 0.998 0.998
gafgyt_udp 0.997 0.997
mirai_ack 0.958 0.959
mirai_scan 0.999 0.999
mirai_syn 0.999 0.999
mirai_udp 0.957 0.959
mirai_udpplain 0.997 0.997
VI-D ROC Curve
The Receiver Operating Characteristics curve demonstrates the connection among sensitivity at several thresholds. As illustrated in fig:6, the curve for the hybrid approach indicating great performance. This suggests that model generates a large number of true positives while maintaining a low rate of false positives. These results presented that the respective hybrid approach is effective in distinguishing between classes like negative and positive.
Figure 7: ROC Curve
VII Comparative Analysis with Similar Architectures
TABLE IV: Comparative Analysis
Class 1D-CNN Precision[9] 1D-CNN Recall[9] 1D-CNN F1-Score[9] CNN-LSTM Precision[9] CNN-LSTM Recall[9] CNN-LSTM F1-Score [9] 1D-CNN-BiLSTM Attention mechanism Precision 1D-CNN-BiLSTM Attention mechanism Recall 1D-CNN-BiLSTM Attention mechanism F1-Score
Benign 1 1 1 1 1 1 1 0.99 0.99
gafgyt_combo 0.93 0.84 0.88 0.93 0.88 0.90 1 0.96 0.98
gafgyt_junk 0.81 0.94 0.87 0.85 0.94 0.89 0.97 1 0.98
gafgyt_scan 1 1 1 1 1 1 1 1 1
gafgyt_udp 0.5 1 0.66 0.55 1 0.66 1 1 1
mirai_ack 0.99 0.99 0.99 1 1 1 0.93 1 0.96
mirai_scan 0.98 1 0.99 0.99 1 0.99 1 1 1
mirai_syn 1 0.99 0.99 1 0.99 0.99 1 1 1
mirai_udp 0.96 0.91 0.93 1 0.93 0.96 1 0.93 0.96
mirai_udpplain 0.91 0.96 0.93 0.93 0.97 0.95 1 1 1
Table IV, demonstrates a contrast between [9] and our model in terms of evaluation metrics. In Table IV, the proposed hybrid deep learning strategy defeated all other models, obtaining superior accuracy than other approaches. Higher Precision and F1-Scores from Comparison Table: Our proposed approach deliberately achieves greater precision and F1-scores among several classes types, demonstrating well accuracy and balance. Perfect Scores in Critical Classes: For critical classes like Bashlite UDP and Mirai UDPplain, the proposed model achieves perfect scores, demonstrating superior performance. Balanced Performance: Even in classes where our model doesn’t achieve perfect scores, it maintains a good balance between precision and recall, obtaining a large output of F1-scores. Overall, the 1D-CNN+BiLSTM+Attention model demonstrates superior performance, particularly in achieving higher precision and F1-scores, making it more reliable and effective for the given classification task.
VIII Conclusion
This paper proposed and evaluated a deep learning model for botnet attack detection, which incorporated an attention process into a hybrid 1D-CNN and Bi-LSTM architecture. The proposed model demonstrated 99% accuracy and close to ideal results for reliability parameters including Matthews Correlation coefficient and Cohen’s kappa coefficient. Moreover, the optimized architecture ensured minimal computation costs, making it perfect for real-time monitoring and protection of IoT devices.